78,088.72J of heat is added to 1000 g of ice at -10 oC. How much ice melts?
Answer #1
78,088.72 = 1000X 2.1 X 10 + m X 334000

solve for m
Answer #2

Heat required to take ice to 0 oC=mcΔT=1000*2.1*10=21,000 J

Remaining heat=77079.1-21000=56079.1 J

Amount of ice that will melt= Remaining heat/Lf= 56079.1/334 =167.9 g
Answer #3
59468.5 = m(335000). m = 177.5g